OpenStudy (anonymous):

evaluate the iterated integral by converting to polar coordinates int_(from -a to a) int_(from 0 to sqr(a -y^2) of (x^2 + y^2) ^3/2 dxdy please help

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

hihi, don't know how to put iint. .whatever, my outer limit from -pi/2 to pi/2 , my inner limit from 0 to sqr (a) the function is r^4 drdtheta.

OpenStudy (anonymous):

\[\huge\iint\r^4drdtheta\]

OpenStudy (anonymous):

I'm going to guess you meant (otherwise it would not be a nice polar integral):  $$\int_{-a}^a\int_0^\sqrt{a^2-y^2} (x^2+y^2)^\frac32\,\mathrm{d}x\mathrm{d}y=\int_{-\frac\pi2}^{\frac\pi2}\int_0^ar^\frac53\,\mathrm{d}r\,\mathrm{d}\theta$$